OpenStudy (anonymous):

A triangle has a leg that measures fifteen and a hypotenuse of twenty-three. Which leg measure would form a right triangle? fifteen thirty-nine square root of three hundred four the square root of seven hundred eighty-five

OpenStudy (anonymous):

@sammixboo @HelpBlahBlahBlah @kewlgeek555 @Destinymasha @d3Xter @SeaTurtle113 @arabpride

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

Know pythagoras?

OpenStudy (anonymous):



OpenStudy (anonymous):

Do you mean a^2+ b^2= c^2?

OpenStudy (anonymous):

\[x^2 = 23^2 - 15^2 \rightarrow x = \sqrt{529-225}\]

OpenStudy (anonymous):

Evaluate the expression. There's your answer :)

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

Use the Pythagoras formula: x^2 = (23)^2 - (15)^2 = 304 x = V304 = ....

OpenStudy (anonymous):

I didn't learn that though...

OpenStudy (anonymous):

It's basically the same as this a^2+ b^2= c^2, but in this case we're looking for a^2. See?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

What is x then?

OpenStudy (anonymous):

So couldn't I use c^2-b^2= a^2?

OpenStudy (anonymous):

And x is a^2

OpenStudy (anonymous):

Exactly, then you'd take square root of both sides to get just a

OpenStudy (anonymous):

c^2= 23^2 b^2= 15

OpenStudy (anonymous):

I substituted x for a

OpenStudy (anonymous):

yes, ... b^2 = 15^2

OpenStudy (anonymous):

23^2- 15^2= 8^2 -------- Now, the square root of 8^2

OpenStudy (anonymous):

So 8*8= 64

OpenStudy (anonymous):

No, evaluate 23^2 before subtracting

OpenStudy (anonymous):

Evaluate?

OpenStudy (anonymous):

Ohhhh nvm

OpenStudy (anonymous):

So, you have 529 - 225

OpenStudy (anonymous):

Can't you do it either way?

OpenStudy (anonymous):

Wait no

OpenStudy (anonymous):

No, you can't! It's against the rules of mathematics.

OpenStudy (anonymous):

I know... sorry I answered myself right after I asked

OpenStudy (anonymous):

So sqrt of 304?

OpenStudy (anonymous):

Good, what is it?

OpenStudy (anonymous):

It doesn't require that... but it's around 17.43

OpenStudy (anonymous):

Okay, I see your options were just left at that point. There's your answer!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Yw :)

OpenStudy (anonymous):

Sorry to be a pain!